OpenStudy (anonymous):

Match the equation with the correct solution. log 7 24 - log 7 (x + 5) = log 7 8

OpenStudy (anonymous):

here are the answer choices 2 -2 5 12 4

OpenStudy (anonymous):

@mathhasproblems @mahmit2012 @imilee @eyust707 can anyone of yall help or find me someone that can help?

OpenStudy (anonymous):

The answer is 5

OpenStudy (anonymous):

how do you know?

OpenStudy (anonymous):

how do you kno @HayleyWhite17

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

this is a similar question to before... the base of the logs are the same... except subtraction indicates division \[\log_{7} 24 - \log_{7} (x + 5) = \log_{7}\frac{24}{(x+5)}\] then you can equate the left and right sides and solve for x \[\log_{7} \frac{24}{x +5} = \log_{7} 8\]

OpenStudy (anonymous):

ohk so where do i go from there?